# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 7.1.2 Detection efficiency
Figure 14 and Table 2 detail the passage and Gabriel cell detection time in experiments. Passage detection time using Delaunay graphs and the direct brute-force traversal strategy are reported. While data in both fixed and varying obstacle sizes are provided, the detection time is similar for the same obstacle number since obstacle size variations do not affect the detection complexity. How Delaunay graphs improve the detection efficiency and how the detection time varies with the obstacle number are of interest. Because of fast Delaunay graph construction and the subsequent linear passage check routine within a small geodesic distance in the graph, significant time is saved compared to directly traversing all obstacle pairs as demonstrated in Figure 14, 14, and Table 2. This improvement becomes more obvious as the obstacle number increases. Up to an order of magnitude speedup is reached, e.g., when 200 obstacles are present. Figure 14 indicates that the passage detection time increasing trend in Delaunay graphs is much less drastic than the direct check method, which has a cubic time complexity in the worst case. Due to the dominance of linearity in $O(m\log m)$ complexity, linearity between the detection time and obstacle number is obvious when using Delaunay graphs.
Figure 16: CPW-PTOPP examples imposed on different types of planners. Obstacles of random shapes and sizes are uniformly distributed and the obstacle numbers in each row are 20, 40, and 60, respectively. (a)-(c) CPW-PTOPP is integrated into shortest path planning. (d)-(f) CPW-PTOPP is integrated into MPW-PTOPP and GPW-PTOPP.
[…]
@@ 7.2 PTOPP results
After passage and Gabriel cell detection, all PTOPP categories and the dependent subroutines are implemented and thoroughly tested. This part reports PTOPP results and particularly focuses on accessible free space optimization performances characterized by different path cost formulations compared to benchmarks. As a new path planning paradigm leveraging passage properties, PTOPP requires new functionalities in planners that have not been realized before, e.g., passage detection and basic primitives in Section 5. Despite many efficient planner implementations available, e.g., the open motion planning library (Sucan et al., 2012), there are no existing interfaces for passage processing, traversal check, and cost configuration to develop PTOPP algorithms based on them readily. PTOPP planners are implemented independently with all subroutines in C++. In particular, PRM and RRT for PTOPP are tested as representative sampling-based methods. PTOPP primitives are integrated as common components invokable by upper-layer planners. Built on these primitives, PTOPP categories are easily configurable in planner instances via specifying cost computation ways. The test environments cover 2D and 3D workspaces. Though the chief objective of PTOPP is to enable large accessible free space along paths, no explicit clearance to obstacles is embedded in the path planning stage. Paths are optimal in the sense of having large accessible free space, not clearance, available over their lengths. But clearance is easily achievable by postprocessing because large clearance is essentially admitted in PTOPP. For instance, raw paths may be locally translated to passage centers to enlarge clearance (Huang et al., 2023, 2024b). Such postprocessing can be done efficiently to make practically usable paths. For consistency, path clearance is not explicitly constrained in the planning stage of PTOPP.
@@ 7.2.1 PTOPP results by categories
Figure 15 presents path planning examples in 2D maps in which RRT for both MPW-PTOPP and GPW-PTOPP are conducted, and shortest path planning acts as the benchmark. The planned paths commonly differ in traversed passages in various planners. In planar environments, the differences essentially correspond to different path homotopy classes. Instead of being the shortest, paths in PTOPP pass selected passages to optimize accessible free space characterized by associated costs. The start and goal positions are placed at the top left and bottom right corners, respectively. 3-GPW-PTOPP is implemented with a weight item common ratio of $100$ in experiments. Obstacles of varying sizes respect uniform or Gaussian distributions for diverse distributions. Environment boundaries are processed as obstacles. Each plan is run until the maximum valid sample number $|V|=5$k, which is sufficient for optimal path convergence in tests. Raw paths may graze obstacles since no postprocessing for path clearance is carried out, but the accessibility of free space is not influenced.
@@
Figure 17: Examples of PTOPP results with different path costs in 3D space. Obstacles are uniformly and randomly distributed in the environment and obstacle numbers are 20, 40, and 60, respectively. (a)-(c) Side views. (d)-(f) Corresponding top views.
The shortest paths (red paths) typically pass through quite narrow passages to achieve the minimum lengths, leading to limited accessible free space alongside. In contrast, paths in PTOPP traverse particular passages to admit large free space. Although both improve the accessibility of free space, MPW-PTOPP (green paths) and GPW-PTOPP (blue paths) behave differently. In general, GPW-PTOPP is more greedy in choosing wide passages. For instance, in Figure 15 and 15, narrow passages enclose the start position and thus are inevitable. After passing them at the beginning part of the path, MPW-PTOPP degrades in the sense that for passages encountered in the future, only those narrower than the previously passed ones are actively avoided. In consequence, some narrow passages are traversed in MPW-PTOPP when they could have been avoided. This problem is alleviated in GPW-PTOPP which optimize multiple confined passages more than the narrowest one. Therefore, GPW-PTOPP paths in Figure 15 and 15 have selected routes enabling significantly larger free spaces alongside. The sorted passage width vector $\mathbf{p}_{\sigma}$ in GPW-PTOPP is guaranteed to be lexically greater than or equal to that in MPW-PTOPP. In many cases, however, two PTOPP planners return paths traversing identical passages, namely homotopic paths in 2D space, like in Figure 15.
Constrained passage width PTOPP is tested by assigning the minimum admissible passage width $\underline{\varepsilon}_{p}$, which represents a ubiquitous demand in robotics to accommodate robots. It is readily integrated into shortest path planning as in (25) or PTOPP via relabeling passage widths as in (26) to penalize traversing unqualified narrow passages. Figure 16 presents CPW-PTOPP results in different setups ($\underline{\varepsilon}_{p}=60$). Figure 16 to 16 are built on shortest path planning. The dashed and solid paths are without and with the constrained passage width (CPW), respectively. CPW-PTOPP finds the shortest path avoiding passages narrower than $\underline{\varepsilon}_{p}$. If there is no feasible path with its minimum passage width $f_{p}(\sigma)>\underline{\varepsilon}_{p}$, CPW-PTOPP will pass the fewest passages narrower than $\underline{\varepsilon}_{p}$ instead of reporting failure as in conventional approaches, e.g, imposition of clearance constraint. Figure 16 to 16 display planned paths in MPW-PTOPP and GPW-PTOPP with CPW. $f_{p}(\sigma)$ is maximized in MPW-PTOPP and GPW-PTOPP even without an admissible threshold $\underline{\varepsilon}_{p}$. Therefore, similar homotopic paths are found in most situations since relabeling passage widths does not affect greedy selection of wide passages, e.g., Figure 16 and 16. However, CPW-PTOPP is still beneficial. Specifically, if $f_{p}(\sigma)\leq\underline{\varepsilon}_{p}$ for all feasible paths, CPW-PTOPP finds the path traversing passages narrower than $\underline{\varepsilon}_{p}$ as few as possible. For instance, there is no feasible path with $f_{p}(\sigma)>\underline{\varepsilon}_{p}$ in Figure 16. Without CPW, the first three items of the sorted passage vector $\mathbf{p}_{\sigma}$ of the planned path are $[59.1,59.3,60.2]^{\mathsf{T}}$. With CPW, items are $[42.3,60.2,63.9]^{\mathsf{T}}$, in which only one entry is smaller than $\underline{\varepsilon}_{p}$ in spite of a smaller $f_{p}(\sigma)$. The overall path cost goes down due to fewer violations of the passage width constraint.
Figure 18: Paths in PTOPP first reach high heights and then go to the target in 3D space. The obstacle number is 60. (a) $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are diagonal ends of the base ground. (b) $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are diagonal ends of the x-y plane at the same height of 200. (c) $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are diagonal ends of the environment.
@@
Figure 19: Examples of CPW-PTOPP results imposed on different types of planners in 3D space. The numbers of random obstacles in each row are 20, 40, and 60, respectively. (a)-(c) CPW-PTOPP is integrated into shortest path planning. (d)-(f) CPW-PTOPP is integrated into MPW-PTOPP and GPW-PTOPP.
@@ 7.2.2 PTOPP results in 3D space
PTOPP algorithms are further tested in 3D environments, the most common workspace in robotics applications. In implementation, the major modifications come from spatial passage detection and compound Gabriel cell decomposition in preprocessing, and the adjusted node positioning procedure in planning. Node projections are positioned in the planar cells on the base for fast passage traversal check. To check if a path traverses midair passages, they are added as virtual boundaries of cells that their projections intersect with. To obtain the passed passages of an edge, both real and virtual cell boundaries are examined. In this way, passage traversal status in 3D space is efficiently computed in PTOPP.
Figure 17 showcases PTOPP results in 3D space returned by RRT. Different planning categories are conducted in the same environment. The 3D environment dimension remains unchanged, and the boundaries are not depicted. The start and goal positions are set as the bottom right and top left corners respectively, with height going from 100 to 200. In addition to displacement on the x-y plane, the height goes up significantly. The maximum valid sample number increases to $|V|=15$k. The shortest paths serve as a benchmark. PTOPP improves the free space accessibility in different ways. Specifically, obstacles of different heights restrict the free space, and the shortest paths usually traverse narrow passages to minimize the path length. MPW-PTOPP and GPW-PTOPP enlarge accessible free space by selecting wider passages to travel. Analogously to the planar case, MPW-PTOPP only optimizes the narrowest passage width. Although accessible free space is enlarged in MPW-PTOPP, GPW-PTOPP maximizes the top k narrowest passages, making it more greedy in optimizing accessible free space. As demonstrated in Figure 17, paths returned by GPW-PTOPP ($\textit{k}=3$) usually have longer lengths in order to pass wider passages.
@@
The paths additionally go up and down to traverse wide passages in 3D space. If $\mathbf{x}_{0}$ or $\mathbf{x}_{g}$ is at a relatively low height, the optimal paths prefer high heights because passages are sparse and wide there. This is exemplified in Figure 18 where at least one of $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ has a low height. The optimal paths in PTOPP tend to go upwards first to reach a high height and then lead to the goal position. By doing this, the paths only go through a few wide passages at high heights, contributing to the minimized path cost. In practice, this strategy may be undesired if certain properties are expected from paths. For instance, robots need to traverse different heights sufficiently for information collection purposes in search tasks or have constrained height climb capabilities. To address this, the sampling strategy can be specified, such as limiting the admissible height range of samples. For this reason, the height range of sampling in Figure 17 is set as $[0,205]^{\mathsf{T}}$ with the upper bound slightly larger than $\mathbf{x}_{g}$ height to prevent the optimal paths from going too high. At last, CPW-PTOPP is also tested in 3D space. As showcased in Figure 19, $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ remain the same as in Figure 17. The impact of a constrained passage width is similar to planar cases. If imposed in shortest path planning, it avoids spatial passages narrower than the admissible width ($\underline{\varepsilon}_{p}=60)$ as much as possible, i.e., passing no such passages or the fewest possible. If imposed in PTOPP, it minimizes the number of passed passages whose widths are smaller than $\underline{\varepsilon}_{p}$.
@@
Table 3: Average path planning time using different planners in 2D environments (ms). Data are collected from 30 trials for each obstacle number. In each trial, planners are executed in the same environment. The total sample number in each plan is $|V|=2$k.
@@
Table 4: Average path planning time using different planners in 3D environments (ms).
@@
7.3 Computational efficiency of PTOPP
At last, the computational efficiency of the proposed PTOPP algorithms is reported. Two major aspects are investigated: PTOPP efficiency compared with benchmark algorithms and previous avenues. The benchmarks include the vanilla shortest path planning and two common approaches to reaching large free space for paths, i.e., planning under a clearance constraint and planning to maximize clearance. Planning under a clearance constraint restricts paths inside the interior of the free space $\mathcal{X}_{free}$. For simplicity, this is done by discarding samples of a clearance smaller than the threshold and assuming edges connecting valid samples lie in the interior. Maximizing clearance is achieved by setting the negative path clearance as the path cost, namely $f_{c}(\sigma)=-\text{clr}(\sigma,\mathcal{X}_{obs})$, where $\text{clr}(\sigma,\mathcal{X}_{obs})$ designates the clearance between $\sigma$ and obstacles. Similarly to the minimum passage width, $f_{c}(\sigma)=-\text{clr}(\sigma,\mathcal{X}_{obs})$ is compatible to be optimizable by sampling-based optimal planners. GPW-PTOPP is run as a representative of PTOPP for its complete use of all developed modules. The compared previous implementation is in (Huang et al., 2024a) that has no sample positioning in Gabriel cells and needs to traverse all passages when checking the passed passages of edges. These kinds of planners are built both on PRM and RRT.
Figure 20: Average path planning time of different planners with different obstacle numbers. Dashed and solid lines correspond to PRM-based and RRT-based algorithms, respectively. For each obstacle number, 30 trials in randomly generated obstacle distributions are conducted.
Figure 21: Average path planning time of different planners under different valid sample numbers. For each valid sample number, five trials in the same environment with 40 obstacles are conducted.
The preprocessing time of passages and cells is negligible compared to the planning phase and is not taken into account. Each planning time is averaged across 30 random obstacle distributions for the same distant setup of $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$. A plan is terminated when the valid sample number reaches $|V|=2$k. Clearance computation involves calculating the smallest distance between samples and obstacles, which can be rapidly performed using R-trees or geometric computations for simple shapes. It is carried out in constrained clearance and maximum clearance planning. A larger clearance constraint incurs more sampling trials to achieve the valid sample number because the feasible sampling space becomes smaller. To be fair, a relatively small clearance constraint is adopted (10 in tests). Another parameter influencing the planning time is the edge length $\eta$ in RRT Steer() subroutine. It decides the range of near neighbors search in Near(). In PTOPP, a longer edge will traverse more cells in PositionNode(). A larger $\eta$ thus means increased computational load. We set $\eta=50$ in 2D tests and $100$ in 3D tests so that they are of similar magnitudes of $r_{PRM^{*}}$ for better comparability.

### v2

@@ 7.1.2 Detection efficiency
Figure 12 and 12 detail the passage and Gabriel cell detection time. Passage detection time using Delaunay graphs and direct brute-force traversal is reported. How Delaunay graphs improve the detection efficiency and how the detection time varies with the obstacle number are of interest. Because of fast Delaunay graph construction and the following linear passage check routine within the graph, significant time is saved compared to directly traversing all obstacle pairs in Figure 12. This improvement becomes more evident when obstacles increase and up to an order of magnitude speedup is reached, e.g., when $m=200$. The passage detection time rising trend in Delaunay graphs is much less drastic than the direct check method, which has cubic time complexity. Due to the dominance of linear complexity in $O(m\log m)$, linearity between detection time and obstacle number is obvious in Delaunay graphs.
Figure 12: Statistics in passage and cell detection. (a), (b) 2D passage and cell numbers under varying obstacle number $m$. (c) Passage numbers when $m=40$ and obstacles vary in size. (d) Passage and cell numbers under varying $m$ in 3D space. (e) Passage detection time based on Delaunay graphs and direct traversal check. (f) Passage and compound cell detection time in 3D space.
Subsequent Gabriel cell detection only performs sorting and DFS of the detected passage graph once. It is thus fast and completes within less than $0.1\text{\,}\mathrm{ms}$ in experiments. In other words, Gabriel cell detection consumes negligible computations compared to passage detection. Passage and cell detection time in 3D space is displayed in Figure 12. The corresponding planar passage detection time is listed as a benchmark. The results show that detecting spatial passages is more time-consuming than planar passages because all intermediate Delaunay graph edges are examined. Although the computational complexity remains, the detection time rises more drastically with obstacle number. Compound cell detection in 3D space is faster than 3D passage detection, but costs much longer time than 2D cells. This time increase originates from processing midair passages. For each midair passage, the base cells its projection intersects with are computed, which accounts for most of the compound cell detection time.
@@ 7.2 PTOPP results
After passage and cell detection, all PTOPP categories are implemented and thoroughly tested. This part reports PTOPP results and particularly focuses on accessible free space optimization results in different path costs. As a new path planning paradigm leveraging passage properties, PTOPP depends on new functionalities that have not been realized before. Despite many efficient planner implementations available such as the open motion planning library (sucan2012open), there are no existing interfaces for passages to develop PTOPP algorithms readily. To address this, planners are implemented independently with all subroutines in C++. PRM and RRT for PTOPP are tested as representatives. Primitives are developed as invokable components and problem categories are easily configurable via specifying cost types. Though the chief objective of PTOPP is to enable large accessible free space along paths, clearance is not embedded explicitly in the planning stage. But it is easily achievable by postprocessing. For instance, raw paths may be locally translated to passage centers to enlarge clearance efficiently (huang2023deformable; huang2024robot). For consistency, path clearance is not explicitly constrained in the planning stage of PTOPP.
@@ 7.2.1 PTOPP results by categories
Figure 13 presents path planning examples in 2D maps. RRT for both MPW-PTOPP and GPW-PTOPP is conducted, and shortest path planning acts as the benchmark. The start $\mathbf{x}_{0}$ and goal $\mathbf{x}_{g}$ are placed at the top left and bottom right corners, respectively. 3-GPW-PTOPP is implemented with a weight item common ratio of $100$. Varying-size obstacles respect uniform or Gaussian distributions, and environment boundaries are processed as obstacles. Each plan runs until the valid sample number |V|=5|V|=5k, which is sufficient for optimal path convergence in tests. Planned paths commonly differ in traversed passages in different planners. PTOPP selects passages to optimize accessible free space characterized by associated costs. In planar maps, such differences essentially infer different path homotopy classes.
@@
Figure 13: PTOPP examples with different cost and environment setups. (a)-(c) Obstacles are uniformly distributed and $m=$ 20, 40, and 60, respectively. (d)-(f) The same numbers of obstacles respect Gaussian distributions.
@@
Figure 14: CPW-PTOPP examples imposed on different types of planners. Obstacles are uniformly distributed. (a)-(c) CPW-PTOPP is integrated into shortest path planning. (d)-(f) CPW-PTOPP is integrated into MPW-PTOPP and GPW-PTOPP.
The shortest paths (red paths) typically pass through quite narrow passages to achieve the minimum lengths, leading to limited accessible free space alongside. In contrast, paths in PTOPP traverse particular passages to admit large free space. Although both improve the accessibility of free space, MPW-PTOPP (green paths) and GPW-PTOPP (blue paths) behave differently. In general, GPW-PTOPP is more greedy in choosing wide passages. For instance, in Figure 13 and 13, narrow passages enclose the start position and thus are inevitable. After passing them at the beginning, MPW-PTOPP degrades in the sense that for passages encountered, only those narrower than the previously passed ones are actively avoided. In consequence, narrow passages are traversed in MPW-PTOPP when they could have been avoided. This problem is alleviated in GPW-PTOPP by optimizing multiple confined passages more than the narrowest one. GPW-PTOPP paths in Figure 13 and 13 select routes enabling significantly larger free spaces alongside. The sorted passage width vector $\mathbf{p}_{\sigma}$ in GPW-PTOPP is guaranteed to be lexically no smaller than that in MPW-PTOPP. In many cases, however, two planners return paths traversing identical passages, namely homotopic paths in 2D space, like in Figure 13.
Figure 15: Examples of PTOPP results with different path costs in 3D space. The environment boundaries are not depicted. Obstacles are uniformly distributed. $m=$ 20, 40, and 60, respectively. (a)-(c) Side views. (d)-(f) Corresponding top views.
Constrained passage width PTOPP is tested by assigning the minimum admissible passage width $\underline{\varepsilon}_{p}$, which represents a ubiquitous demand to accommodate robots. It is readily integrated into shortest path planning as in (25) or PTOPP via relabeling passage widths as in (26) to penalize traversing unqualified narrow passages. Figure 14 to 14 present CPW-PTOPP results built on shortest path planning ($\underline{\varepsilon}_{p}=60$). Dashed and solid paths are without and with the constrained passage width (CPW), respectively. CPW-PTOPP finds the shortest path avoiding passages narrower than $\underline{\varepsilon}_{p}$. If there is no feasible path with its minimum passage width $f_{p}(\sigma)>\underline{\varepsilon}_{p}$, CPW-PTOPP will pass the fewest passages narrower than $\underline{\varepsilon}_{p}$. Figure 14 to 14 display planned paths in MPW-PTOPP and GPW-PTOPP with CPW. $f_{p}(\sigma)$ is maximized in MPW-PTOPP and GPW-PTOPP even without an admissible threshold $\underline{\varepsilon}_{p}$. Therefore, homotopic paths are found in most situations since relabeling passage widths does not affect greedy selection of wide passages, e.g., Figure 14 and 14. However, CPW-PTOPP is still beneficial. Specifically, if $f_{p}(\sigma)\leq\underline{\varepsilon}_{p}$ for all feasible $\sigma$, CPW-PTOPP finds the path traversing passages narrower than $\underline{\varepsilon}_{p}$ as few as possible. For instance, there is no feasible path with $f_{p}(\sigma)>\underline{\varepsilon}_{p}$ in Figure 14. Without CPW, the first three items in $\mathbf{p}_{\sigma}$ of the planned path are $59.1,59.3,60.2]^{\mathsf{T}}$. With CPW, items are $42.3,60.2,63.9]^{\mathsf{T}}$, in which only one entry is smaller than $\underline{\varepsilon}_{p}$ in spite of a smaller $f_{p}(\sigma)$. The overall path cost decreases due to fewer violations of the passage width constraint.
@@ 7.2.2 PTOPP results in 3D space
PTOPP algorithms are further tested in 3D space. In implementation, major modifications arise from spatial passage and compound Gabriel cell detection in preprocessing and the adjusted node positioning procedure in planning. Midair passages are added as virtual boundaries of planar cells that their projections intersect with. Node projections are positioned in the planar cells for fast passage traversal check. To obtain the passed passages of an edge, both real and virtual cell boundaries are examined. Figure 15 exhibits PTOPP results in 3D space returned by RRT. $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are set as the bottom right and top left corners respectively with height going from 100 to 200. The maximum valid sample number increases to |V|=15|V|=15k. PTOPP improves the free space accessibility in different ways. Obstacles of different heights restrict the free space. MPW-PTOPP and GPW-PTOPP enlarge accessible free space by selecting wider passages to travel. MPW-PTOPP only optimizes the narrowest passage width. GPW-PTOPP maximizes top k narrowest passages usually at the expense of longer path lengths, making it more greedy in optimizing accessible free space.
@@
Paths will go up and down to traverse wide passages in 3D space. The optimal paths prefer large heights because passages are sparse and wide there. This is exemplified in Figure 16 where $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are low. Optimal paths in PTOPP tend to go upwards first to reach a large height and then lead to the goal. By doing this, the paths only go through a few wide passages at large heights, contributing to a minimized path cost. In practice, this strategy may be undesired if certain path properties are expected. For instance, robots need to traverse different heights for information collection purposes in search tasks. To address this, the sampling strategy can be specified by limiting the admissible height range of samples. The sampled height range in Figure 15 is $0,205]^{\mathsf{T}}$ with the upper bound slightly larger than $\mathbf{x}_{g}$ height to prevent the optimal paths from going too high. Then, CPW-PTOPP is also tested in 3D space as in Figure 17. The impact of a constrained passage width is similar to planar cases. If imposed in shortest path planning, it avoids spatial passages narrower than the admissible width ($\underline{\varepsilon}_{p}=60)$ as much as possible while being the shortest. If imposed in PTOPP, it minimizes the number of passed passages whose widths are smaller than $\underline{\varepsilon}_{p}$.
Figure 16: $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are diagonal ends of the ground. Paths in PTOPP first reach high and then go to the target.
Figure 17: Examples of CPW-PTOPP results imposed on different planner types in 3D space. $\mathbf{x}_{0},\mathbf{x}_{g}$ are the same as Figure 15. (a)-(c) CPW-PTOPP is integrated into shortest path planning. (d)-(f) CPW-PTOPP is integrated into MPW-PTOPP and GPW-PTOPP.
7.3 Computational efficiency of PTOPP
At last, the computational efficiency of the proposed PTOPP algorithms is reported. Two major aspects are investigated: PTOPP efficiency compared with benchmark algorithms and previous avenues. The benchmarks include the shortest path planning and two common approaches to reaching large free space for paths, i.e., planning under a clearance constraint and to maximize clearance. Planning under a clearance constraint restricts paths inside the interior of $\mathcal{X}_{free}$. For simplicity, this is done by discarding samples of a clearance smaller than the threshold and assuming edges connecting valid samples lie in the interior. Maximizing clearance is achieved by setting the negative path clearance as a compatible path cost, namely $f_{c}(\sigma)=-\text{clr}(\sigma,\mathcal{X}_{obs})$, where $\text{clr}(\sigma,\mathcal{X}_{obs})$ designates the clearance between $\sigma$ and obstacles. GPW-PTOPP is run as a representative of PTOPP for the full use of all developed modules. The compared previous implementation is in (huang2024homotopic) that has no sample positioning in Gabriel cells and needs to traverse all passages when checking the passed passages of edges. Planners are built both on PRM and RRT.
The preprocessing time of passages and cells is negligible compared to the planning phase and is not taken into account. Each planning time is averaged across 30 random obstacle distributions for the same distant setup of $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$. A plan is terminated when the valid sample number reaches |V|=2|V|=2k in uniform sampling. Clearance calculates the smallest distance between samples and obstacles, which can be rapidly performed using R-trees or geometric computations for simple shapes. A larger clearance constraint incurs more sampling trials to achieve $|V|$ because the feasible sampling space is reduced. To be fair, a relatively small clearance constraint is adopted (10 in tests). Another parameter influencing the planning time is the edge length $\eta$ in RRT Steer() subroutine. It decides the range of near neighbors search. In PTOPP, a longer edge will traverse more cells in PositionNode(). A larger $\eta$ thus means increased computational load. We set $\eta=50$ in 2D tests and $100$ in 3D tests so that they are of similar magnitudes of $r_{PRM^{*}}$ for better comparability.
@@ 7.3 Computational efficiency of PTOPP
Table 1: Planners’ average path planning time in 2D environments (ms). Data are collected from 30 trials for each obstacle number. Each trial is executed in the same environment. The total sample number in each plan is |V|=2|V|=2k.
@@
Table 2: Average path planning time in 3D environments (ms).